How to solve (x-7/5)(x+5/7)?
Answer by stanbon(75887)

How to solve (x-7/5)(x+5/7)?
----
General Form: (a-b)(a+b) = a^2 - b^2
------
Your Problem:
(x-7/5)(x+5/7)
= x^2 - (25/49)
===================
Cheers,
Stan H.